# DateLogicQA: Benchmarking Temporal Biases in Large Language Models

### v1

@@ Abstract
This paper introduces DateLogicQA, a benchmark with 190 questions covering diverse date formats, temporal contexts, and reasoning types. We propose the Semantic Integrity Metric to assess tokenization quality and analyse two biases: Representation-Level Bias, affecting embeddings, and Logical-Level Bias, influencing reasoning outputs. Our findings provide a comprehensive evaluation of LLMs’ capabilities and limitations in temporal reasoning, highlighting key challenges in handling temporal data accurately. The GitHub repository for our work is available at https://github.com/gagan3012/EAIS-Temporal-Bias 111ϕ Denotes equal contribution
@@ 1 Introduction
This paper makes several contributions to understanding temporal biases in LLMs. (1) We introduce DateLogicQA, a dataset of $190$ curated questions for evaluating temporal reasoning across various date formats, contexts (past, present, future), and reasoning types (commonsense, factual, conceptual, numerical). (2) We propose the Semantic Integrity Metric to assess tokenization quality, punishing unnecessary splits and excessive token counts. (3) We conduct human evaluations of model responses to analyse tokenization accuracy and reasoning quality, providing insights beyond automated metrics. (4) A thorough bias evaluation examines representation and logical-level biases using embeddings and outputs to investigate LLMs’ treatment of temporal references.
We have organised the paper as follows: Section 2 reviews related works, summarising the impact of tokenization on LLM performance and past temporal reasoning approaches. Section 3 details the creation of the DateLogicQA dataset, including its design principles and examples. Section 4 outlines methods for evaluating tokenization, temporal reasoning, and biases. Section 5 presents experiment results, followed by a discussion of findings and bias mitigation in Section 6. Lastly, Section 7 summarises our contributions.
@@ 4 Methodology
The study proposes three interests to investigate temporal bias in models: tokenization process, temporal task capability, and internal computation across different LLMs.
4.1 Semantic Integrity
This experiment targets the tokenization process in different LLMs to identify how it influences the semantic interpretation of dates when presented in different formats. We specifically focus on the Semantic Integrity Metric, which measures the extent to which the original semantic meaning of a date is preserved after tokenization. Several key highlights are observed, such as how a single date input is presented after being tokenized, the extent of semantic preservation, and the ability to generalise across different date representations. These findings provide valuable insights into the tokenization process and its impact on temporal reasoning in LLMs.
Semantic integrity evaluates how well the tokenized date output maintains its original meaning and structure. The semantic integrity score ranges from 0 to 1. A higher score - closer to 1 - indicates that the date segmentation is nearly accurate, better preserving the intended structure and information. In contrast, a score closer to 0 indicates an inadequately tokenized date structure. The formula for calculating semantic integrity is as follows:
$\text{SI}=\max\left(0,\min\left(1,1-P-S-T-R\right)\right)$
Unnecessary Splitting of Components (P)
: If a date is tokenized into parts that do not correspond to the ideal format, a penalty of 0.1 is applied.
Preservation of Separators (S)
: A penalty of 0.1 is applied when separators are lost during tokenization, reflecting incorrect date parsing (e.g., tokenizing %Y-%m-%d as %Y%m%d).
Token Count (T)
This penalty design aims to penalise token outputs longer than necessary, such as %Y-%m-%d. If any tokenized output exceeds five splits, we apply a penalty for excessive fragmentation. A higher token count often indicates a loss of semantic information, hindering the model’s ability to interpret temporal values.
Similarity with Baseline (R)
This metric uses cosine similarity to measure how much a tokenized output deviates from a baseline reference of date tokens. Tokens with greater deviation incur higher penalties, while those resembling the baseline receive lower penalties.
@@ 4.2 Human-Led Temporal Bias Assessment
4.3 Understanding Temporal Bias
We investigate potential biases in the internal embedding space and softmax computations of large language models (LLMs) when processing texts with different temporal references, such as past, present, and future contexts. Temporal biases in LLMs fall into two main types: Representation-Level Temporal Bias indicates significant differences in internal embeddings across time references, revealing inconsistencies in encoding semantic information. In contrast, Logical-Level Temporal Bias occurs when output probabilities vary for semantically identical inputs due to changes in temporal references.
We established a controlled experimental framework to quantify these biases, analysing embeddings and softmax outputs across three temporal categories: past, present, and future. We measured representation-level biases using cosine similarity between averaged embeddings, with lower similarity indicating greater divergence. We assessed logical-level biases using KL divergence for softmax distributions, where higher divergence reflects substantial probability differences. Additionally, we examined sensitivity to seven date formats from Table 2.
We conducted experiments using Llama 3.2 3B Dubey et al. (2024). For each prompt, we extracted two key outputs: internal embeddings, averaged from all the hidden layers of hidden states, and softmax probabilities, denoting the output distribution over the vocabulary. We analysed temporal biases by comparing these outputs across different references. Simultaneously, we assessed format sensitivity within each reference to determine how tokenization and variations in date formats affect model behaviour.
@@ 5 Results
5.1 Impact of tokenizers
Model SI TC PC PS
Baseline 1.00 4.30 $\checkmark$ $\checkmark$
OLMoE 0.77 5.08 $\approx$ $\checkmark$
OLMo 0.77 5.08 $\approx$ $\checkmark$
Davinci-003 0.75 5.17 $\times$ $\checkmark$
Llama 3 0.74 4.98 $\times$ $\checkmark$
GPT-3.5 0.74 4.98 $\times$ $\checkmark$
GPT-4 0.74 4.98 $\times$ $\checkmark$
GPT-4o 0.74 4.98 $\times$ $\checkmark$
Qwen 0.42 9.30 $\times$ $\checkmark$
Cohere 0.42 9.30 $\times$ $\checkmark$
Gemma 0.42 9.30 $\times$ $\checkmark$
DeepSeek 0.42 9.30 $\times$ $\checkmark$
Llama 2 0.37 10.30 $\times$ $\checkmark$
Mistral 0.37 10.30 $\times$ $\checkmark$
Phi 3.5 0.37 10.30 $\times$ $\checkmark$
Llama 1 0.37 10.30 $\times$ $\checkmark$
Table 3: Performance comparison of various models on semantic integrity, token count, and preservation of components and separators.
Figure 3: Temporal impact on semantic integrity
@@ 5.1 Impact of tokenizers
Figure 6: Representation level Temporal Bias Analysis using LLama 3.2 3B
Figure 7: Logical level Temporal Bias Analysis using LLama 3.2 3B
In Table 3, we present the Semantic Integrity scores with respect to token count (TC), component preservation (PC) and separators (PS) for different tokenizers used by the tested models. Notably, newer released models tend to have a higher average semantic integrity score at $\approx 0.7$ while token counts maintained close to the frequency of $\approx 5$.
As described in Table 2, most of the tested date formats are composed of 5 components (e.g., DD-MM-YY), with some instances of a 3-component structure like the Julian date format (YYYY/DD). As a result, the baseline reference has an average token count of $4.30$ rather than the ideal $5$.
While many tokenizers struggled to consistently preserve the components, often producing tokenized outputs that deviated from the ideal, the tokenizers used by the OLMoE Muennighoff et al. (2024) and OLMo Groeneveld et al. (2024) models handled date inputs correctly on certain occasions.
Token Count and Semantic Integrity
From Table 3, we identify an indirect relationship between the token count and the Semantic Integrity score, specifically at lower Semantic Integrity often corresponds to more token splits, suggesting that excessive and inefficient token splitting could badly impact the interpretability of the original input. Figure 8 further establishes this observation by showing that the Semantic Integrity scores tend to fall near the lower end in the area of higher token count.
Shared Characteristics Between Tokenizers
Results produced by some tokenizers have been identical, as observed in Table 3. For example, the tokenizers that Llama 1 Touvron et al. (2023a) and Mistral Jiang et al. (2023) use produce similar Semantic Integrity scores, a trend also observed in other models. Therefore, we investigated this further by evaluating the output from the tokenizers used by all models, presenting the results in Table 4. This finding seems accurate, as we identified similar tokenized date outputs from the tokenizers used in certain models. For instance, Macijauskas (2024) has also concluded a similar finding for the tokenizer used by GPT-4 OpenAI et al. (2023) and Llama 3 Dubey et al. (2024), which both produced similar outputs during his experiment. From our observation, this helps justify the idea of shared tokenizers between several models.
Performance in Different Temporal References
In addition, Figure 3 reveals that certain tokenizers exhibit a temporal bias, resulting in varying scores produced in different timelines. Some tokenizers yield better Semantic Integrity scores for dates closer to the present (1900s-2100s). Although this is not consistently true across all tokenizers, this pattern highlights a potential bias in a tokenizer, with some favouring more recent dates over others. Slight variations in Semantic Integrity scores are particularly noticeable between the 1700s and 1900s for the tokenizers from Gemma Team et al. (2024) and Phi 3.5 Abdin et al. (2024) model. In contrast, the tokenizers used by OLMoE Muennighoff et al. (2024) and OLMo Groeneveld et al. (2024) exhibited more significant score fluctuations in more recent years.
@@ 5.2 Temporal Reasoning Analysis
The evaluation of 12 language models, accessed through Hugging Face and OpenAI APIs, provided a comprehensive overview of their performance on temporal reasoning tasks. Small models like Llama-3.2-3B Dubey et al. (2024) and Phi-3.5-mini Abdin et al. (2024) gave bad performances, with 58% and 66% incorrect answers, respectively. Due to their restricted processing and resources, these models performed poorly in tokenization and reasoning. Mid-sized models, including Mistral-7B Jiang et al. (2023), Llama-3-8B Dubey et al. (2024), and Llama-2-7B Touvron et al. (2023b), demonstrated a more moderate improvement. They had trouble with complex reasoning problems, although they were able to improve their tokenization accuracy. Larger models, including Llama-3-70B Dubey et al. (2024), Qwen2.5-72B Yang et al. (2024), and Command R+ Cohere (2024), were more robust in their performance, especially in date interpretation and logical reasoning. However, there were inconsistencies in specific formats. Proprietary models, including GPT-3.5 Brown et al. (2020), GPT-4-turbo OpenAI et al. (2023), GPT-4o, and GPT-4o-mini OpenAI et al. (2024) outperformed all the rest, with GPT-4-turbo leading on correct responses with 63% and the lowest rate of incorrect answers at 16%. These results emphasise that model size, architecture, and diversity of pretraining data all bear on performance related to temporal reasoning tasks.
@@
5.3 Temporal Sensitivity Analysis
We analyse temporal biases and format sensitivity by examining the embeddings and softmax outputs of the model for prompts across three temporal categories — past, present, and future — and multiple date formats, as shown in Table 2. We organise the findings into representation-level bias, logical-level bias, and format sensitivity, and we show the results in  Figure 6 and Figure 7.
Representation-Level Bias
We evaluated representation-level bias by calculating the cosine similarity between the averaged embeddings for prompts across the three temporal references. The leftmost heatmap in Figure 6 illustrates these similarities.
The embeddings for past and present references exhibit no measurable similarity (0.00), emphasising that the model encodes historical and contemporary contexts with distinct semantic structures. However, the moderate similarity between future and present suggests some shared contextual features between these categories while maintaining semantic differentiation. The moderate similarity between past and future indicates that these categories share overlapping contextual features while remaining semantically distinct. This further implies that the model is somewhat confused regarding futuristic references, which may frequently misattribute to a different temporal category, likely reflecting the training data distribution.
Logical-Level Bias
We assessed logical-level bias by measuring the Kullback-Leibler (KL) divergence between softmax outputs for prompts across temporal references. The leftmost heatmap in  Figure 7 illustrates these divergences—prompts referencing the present exhibit the lowest divergence, indicating stable and consistent output probabilities. However, significant divergence is observed between past-present and future-present comparisons, highlighting the model’s reliance on different priors when predicting tokens for noncontemporary contexts.
The moderate divergence between past and future outputs suggests that the model differentiates between these temporal categories while leveraging some shared contextual grounding. The distinct KL divergences for non-present prompts indicate a logical-level bias, where the model’s probabilistic outputs are sensitive to the temporal context, even when the semantic content of the prompts remains equivalent.
5.4 Format Sensitivity Analysis
Figures 6 and 7 (second to fourth columns) show the model’s sensitivity to variations in date formats for each temporal reference. Both embeddings and softmax outputs reveal notable patterns of variability across formats.
Representation-Level Bias
The cosine similarity heatmaps in Figure 6 indicate that date formats with standard separators (e.g., %Y-%m-%d) yield higher consistency, particularly for present references. Non-standard formats (e.g., %Y%m%d, %d%m%Y) result in lower similarity, especially for past and future prompts. The future category exhibits the highest variability in embeddings across formats, suggesting that futuristic contexts rely more on consistent input structures. In contrast, embeddings for present references remain robust across formats, likely due to the dominance of contemporary contexts in the training data.
Logical-Level Bias
The KL divergence heatmaps in Figure 7 reflect similar trends. Standardised data formats (e.g., %Y-%m-%d) produce more stable predictions, while non-standard formats (e.g., %d%m%Y) introduce higher variability. This sensitivity manifests most prominently in future references, where the KL divergence values are consistently higher, indicating that the model’s predictions have increased uncertainty. In contrast, present references remain relatively stable, reinforcing the model’s preference for standardised inputs and contemporary contexts.
The results highlight two fundamental temporal biases in the model. First, representation-level biases reveal that the model encodes temporal contexts with distinct semantic structures, likely shaped by training data distribution. Second, logical-level biases indicate inconsistencies in output probabilities across temporal references, underscoring the challenges of achieving temporal generalisation. Furthermore, the heightened sensitivity to non-standard date formats underscores the importance of input standardisation for ensuring consistent model behaviour in temporal reasoning tasks.
@@ Appendix A Appendix
Appendix B Contributions
Figure 9: Contributions Chart
Table 5: Model Performance on DateLogicQA

### v2

@@ Abstract
We introduce DateLogicQA, a human curated benchmark of 190 questions specifically designed to understand temporal bias in Large Language Models (LLMs). Covering seven date formats across past, present, and future contexts, DateLogicQA examines four reasoning types: commonsense, factual, conceptual, and numerical. Through human-led evaluations of 12 state-of-the-art LLMs, we identify Representation-Level Bias, arising from suboptimal embeddings that distort date semantics, and Logical-Level Bias, manifesting when correct date tokens yield flawed temporal reasoning. Our findings underscore persistent challenges in handling various date formats and temporal contexts, revealing the need for more robust pretraining data, targeted post-training methods, and precise tokenization strategies. By illuminating these biases, we provide actionable insights to guide the development of LLMs for accurate temporal reasoning across diverse real-world applications.
DateLogicQA: Benchmarking Temporal Biases in Large Language Models
Gagan Bhatia ϕ, MingZe Tang ϕ, Cristina Mahanta ϕ, Madiha Kazi ϕ University of Aberdeen {g.bhatia.24,m.tang.24,c.mahanta.24,m.kazi.24}@abdn.ac.uk
@@ 1 Introduction
This paper makes two significant contributions to understanding temporal biases in LLMs. (1) We introduce DateLogicQA, a dataset of $190190$ curated questions for evaluating temporal reasoning across various date formats, contexts (past, present, future), and reasoning types (commonsense, factual, conceptual, numerical). (2) We conduct human evaluations of model responses to analyse tokenization accuracy and reasoning quality, providing insights beyond automated metrics.
We have organised the paper as follows: Section 2 reviews related works, summarising the impact of tokenization on LLM performance and past temporal reasoning approaches. Section 3 details the creation of the DateLogicQA dataset, including its design principles and examples. Section 4 outlines methods for temporal reasoning, and biases. Section 5 presents experiment results, followed by a discussion of findings and bias mitigation in Section 6. Lastly, Section 7 summarises our contributions.
@@ 5.1 Temporal Reasoning Analysis
The evaluation of 12 language models, accessed through Hugging Face and OpenAI APIs, provided a comprehensive overview of their performance on temporal reasoning tasks. Small models like Llama-3.2-3B Dubey et al. (2024) and Phi-3.5-mini Abdin et al. (2024) gave bad performances, with 58% and 66% incorrect answers, respectively. Due to their restricted processing and resources, these models performed poorly in tokenization and reasoning. Mid-sized models, including Mistral-7B Jiang et al. (2023), Llama-3-8B Dubey et al. (2024), and Llama-2-7B Touvron et al. (2023), demonstrated a more moderate improvement. They had trouble with complex reasoning problems, although they were able to improve their tokenization accuracy. Larger models, including Llama-3-70B Dubey et al. (2024), Qwen2.5-72B Yang et al. (2024), and Command R+ Cohere (2024), were more robust in their performance, especially in date interpretation and logical reasoning. However, there were inconsistencies in specific formats. Proprietary models, including GPT-3.5 Brown et al. (2020), GPT-4-turbo OpenAI et al. (2023), GPT-4o, and GPT-4o-mini OpenAI et al. (2024) outperformed all the rest, with GPT-4-turbo leading on correct responses with 63% and the lowest rate of incorrect answers at 16%. These results emphasise that model size, architecture, and diversity of pretraining data all bear on performance related to temporal reasoning tasks.